OpenStudy (anonymous):

completing the square x^2 +6x -3 =0

OpenStudy (anonymous):

(x+3)^2 - 9 -3 = 0 (x+3)^2 = 12 x+3 = +- sqrt12 x= - 3 + sqrt12 or -3 - sqrt12

OpenStudy (anonymous):

-3 + 3.464 or -3 - 3.464 = 0.464 or -6.464

OpenStudy (lalaly):

\[\large{(x+\frac{6}{2})^2-(\frac{6}{2})^2-3=0}\]\[\large{(x+3)^2-(3)^2-3=0}\]\[\large{(x+3)^2-9-3=0}\]\[\large{(x+3)^2=12}\]\[\large{x+3= \pm \sqrt{12}}\] \[\large{x=\sqrt{12}-3=2\sqrt3-3}\]\[\large{x=-\sqrt{12}-3}=-2\sqrt3-3\]